OpenStudy (anonymous):

Convert z=6cis(π) to rectangular form

OpenStudy (tkhunny):

What does \(cis(\pi)\) mean? Expand that to its full expression.

OpenStudy (anonymous):

I don't know its just one of the problems i have to do for an online class

OpenStudy (anonymous):

Since \(\large \mathrm{cis}(\theta) = \cos\theta+i\sin\theta\), we have that \(\large 6\mathrm{cis}(\pi) =6(\cos\pi + i\sin \pi) = \ldots\) Can you take things from here? :-)

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

z=-6 Thanx! @ChristopherToni

OpenStudy (anonymous):

That's correct. I hope that clarifies things!